OpenStudy (anonymous):

can anyone solve x^2+x+1=0---quadratic equation by completeting the square

OpenStudy (ash2326):

We have \[x^2+x+1=0\] We know that \[(x+a)^2=x^2+2ax+a^2\] here coefficient of x=1 so \[2a=1\] \[a=\frac{1}{2}\] so we need to add and subtract \(\frac{1}{2^2}\) to complete square \[x^2+x+\frac{1}{2^2}-\frac{1}{2^2}+1=0\] we get \[(x+\frac{1}{2})^2 +(\frac{\sqrt{3}}{2})^2=0\] now we get \[(x+\frac{1}{2})^2 =-(\frac{\sqrt{3}}{2})^2\] we know \(\sqrt{-1}=i\) so we get \[(x+\frac{1}{2})=\pm i\frac{\sqrt{3}}{2}\] we get \[x=\frac{-1}{2}\pm i\frac{\sqrt{3}}{2}\] or \[x=\frac{-1\pm i\sqrt{3}}{2}\]

OpenStudy (anonymous):

wow. thanks for be so detailed.

OpenStudy (ash2326):

Welcome, did you understand?

OpenStudy (anonymous):

I understand the formula of needing to find the (b) of 2ab but I got lost in the ^3/2 part

OpenStudy (anonymous):

once i rework the problem I should be able to understand your answer

OpenStudy (ash2326):

Yeah we get \[-\frac{1}{4}+1= \frac{3}{4}= {(\frac{\sqrt 3}{2})}^2\]

OpenStudy (anonymous):

o i c now. that makes sense!